OpenStudy (anonymous):

Eliminate the "z" variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1 A. 4x - 3y = -14 B. 4x + y = -14 C. 6x + 5y = 13 D. 6x - 5y = -13

OpenStudy (whpalmer4):

In the first two equations, the coefficient of z is 2. In the last equation, it is -1. If you double the last equation, it will be -2, and adding them together will cause the elimination of z. After doubling the final equation, line them up with matching terms in columns and add straight down to get your answer.

OpenStudy (whpalmer4):

2x+3y+4z = 6 3x+4y-4z = 10 -------------- 5x+7y+0z=16 or 5x+7y=16 (example only, not the real problem)

OpenStudy (anonymous):

Ok, is this the correct answer? B. 4x + Y =-14

OpenStudy (whpalmer4):

Uh, no. Can you show me your work?